### v1

# Quantum-Adversary-Resilient Evidence Structures and Migration Strategies for Regulated AI Audit Trails

Abstract
Constant-size cryptographic evidence records are increasingly used to build audit trails for regulated AI workloads in clinical, pharmaceutical, and financial settings, where each execution is summarized by a compact, verifiable record of code identity, model version, data digests, and platform measurements. Existing instantiations, however, typically rely on classical signature schemes whose long-term security is threatened by quantum-capable adversaries. In this paper we formalize security notions for evidence structures in the presence of quantum adversaries and study post-quantum (PQ) instantiations and migration strategies for deployed audit logs. We recall an abstraction of constant-size evidence structures and introduce game-based definitions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, capturing the inability of a quantum adversary to forge, equivocate, or rebind evidence items. We then analyze a hash-and-sign instantiation in the quantum random-oracle model (QROM), assuming an existentially unforgeable PQ signature scheme against quantum adversaries, and show that the resulting evidence structure satisfies these notions under standard assumptions. Building on this, we present three migration patterns for existing evidence logs—hybrid signatures, re-signing of legacy evidence, and Merkle-root anchoring—and analyze their security, storage, and computational trade-offs. A case study based on an industrial constant-size evidence platform for regulated AI at Codebat Technologies Inc. suggests that quantum-safe audit trails are achievable with moderate overhead and that systematic migration can significantly extend the evidentiary lifetime of existing deployments.
1 Introduction
Audit trails play a central role in regulated environments where automated computation affects safety- or compliance-critical decisions. In domains such as clinical AI, pharmaceutical development, and high-stakes financial analytics, regulators require that organizations be able to explain and justify how a particular result was produced: which code and model were used, which data were processed, under which policies, on which platform, and at what time. Recent system designs address these needs by representing each execution as a constant-size cryptographic evidence item that compactly captures the essential provenance and integrity information and can be stored and verified at scale (e.g., [kao2025constant, securelogging, transparency]), and recent work has begun to study quantum-resilient designs for distributed log storage itself [somarapu2022qrlogs].
Today, many such systems instantiate their evidence structures with classical primitives such as ECDSA or RSA signatures and SHA-2 family hash functions. These choices are well understood and adequate against classical adversaries, but the long-term evidentiary value of audit logs is threatened by the emergence of quantum computers. A quantum-capable adversary may be able to harvest classical evidence now and forge or equivocate about past executions later, once large-scale quantum computation becomes available. Even for organizations that are not yet deploying quantum hardware, regulators increasingly expect plans for quantum-safe logging and evidence retention.
Despite a growing body of work on post-quantum cryptography (PQC) and secure logging, there is relatively little treatment of evidence structures as a first-class cryptographic object in a quantum-adversary setting. In particular, existing designs do not usually:
  • formalize audit-trail security notions for evidence structures against quantum adversaries;
  • analyze concrete post-quantum instantiations in a standard model such as the quantum random-oracle model (QROM);
  • provide systematic migration strategies for already-deployed evidence logs that must remain verifiable for a decade or longer.
In this work we address these gaps. We build on an abstraction of constant-size evidence structures introduced in prior work [kao2025constant], which models a system where each event in a workflow is associated with a fixed-length evidence record produced by a Generate algorithm and consumed by a Verify algorithm. Our focus here is on extending this abstraction to a quantum-adversary setting, providing both formal definitions and practically oriented migration strategies.
Contributions.
Our main contributions are:
  1. 1.
    Quantum-adversary security notions. We define a system model where adversaries may obtain quantum access to hash oracles and may attempt harvest-now/forge-later attacks against evidence logs. In this model we introduce game-based security notions for evidence structures, including Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding.
  2. 2.
    Post-quantum instantiations in the QROM. We analyze a hash-and-sign instantiation of constant-size evidence structures in the quantum random-oracle model, assuming an underlying post-quantum signature scheme that is existentially unforgeable under chosen-message attacks (EUF-CMA) against quantum adversaries. We prove that under standard assumptions the instantiated scheme satisfies our quantum-adversary security notions.
  3. 3.
    Migration patterns for deployed logs. We study three migration patterns for deployed evidence systems: (i) hybrid classical+PQ signatures for new evidence, (ii) re-signing legacy evidence with PQ signatures in a trusted environment, and (iii) Merkle-root anchoring of legacy batches. For each pattern we discuss trust assumptions, security guarantees, and storage and computational overhead.
  4. 4.
    Industrial case study. We outline a case study using a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc. We discuss how the proposed migration patterns map to realistic deployment constraints and estimate the cost of securing large evidence corpora.
The rest of the paper is organized as follows. Section 2 recalls the abstraction of constant-size evidence structures and reviews quantum-adversary and PQ primitive models. Section 3 introduces our quantum-adversary security notions. Section 4 presents and analyzes a post-quantum instantiation. Section 5 develops migration patterns for deployed evidence logs. Section 6 discusses an industrial case study. Section 7 reviews related work, and Sections 89 conclude with open problems and future directions.
Relation to prior work.
Compared to classical secure logging and transparency systems (e.g., [securelogging, forwardsecurelogging, haberstornetta1991, merkle1989certified, ct, transparency]), our work treats constant-size evidence structures as the primary cryptographic object and explicitly analyzes their security against quantum adversaries in the QROM. Existing PQC work on logging and auditing (e.g., [somarapu2022qrlogs, khan2025pqblockchainaudit, yusuf2025qrlogintegrity]) typically either focuses on blockchain-based storage or on specific application domains. In contrast, we provide (i) generic quantum-adversary security notions for evidence structures, and (ii) migration patterns that are designed to retrofit already-deployed constant-size evidence logs in regulated AI systems.
2 Preliminaries
2.1 Constant-Size Evidence Structures
We briefly recall the abstraction of constant-size evidence structures introduced in [kao2025constant], which we treat here as a black-box building block.
Consider a workflow that generates a sequence of events $e_{1},e_{2},\dots,e_{T}$, where each event $e_{i}$ corresponds to a concrete execution step such as a model inference, a batch data transformation, or an approval action. For each event $e_{i}$ the system produces an evidence record $\mathsf{Ev}_{i}\in\{0,1\}^{\ell}$ of fixed length $\ell$ bits. Informally, $\mathsf{Ev}_{i}$ encodes:
  • identifiers and digests of the code, model, and configuration used;
  • digests of the input and output artifacts;
  • platform measurements (e.g., TEE measurement), timestamps, and policy identifiers;
  • a cryptographic authenticator binding these fields together.
Formally, a constant-size evidence structure is defined by algorithms
$\mathcal{E}=(\mathsf{Setup},\mathsf{KeyGen},\mathsf{Generate},\mathsf{Verify},\mathsf{Link})$
with the following interfaces:
  • $(\mathsf{pp})\leftarrow\mathsf{Setup}(1^{\lambda})$: generates public parameters.
  • $(\mathsf{sk},\mathsf{pk})\leftarrow\mathsf{KeyGen}(\mathsf{pp})$: generates a signing key pair (notation here covers the integrity mechanism; more general forms are possible).
  • $\mathsf{Ev}\leftarrow\mathsf{Generate}(\mathsf{pp},\mathsf{sk},e)$: given an event description $e$, outputs a constant-size evidence record $\mathsf{Ev}$.
  • $b\leftarrow\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e,\mathsf{Ev})$: returns $1$ if $\mathsf{Ev}$ is a valid evidence record for event $e$, and 0 otherwise.
  • $\pi\leftarrow\mathsf{Link}(\mathsf{pp},\mathsf{Ev}_{1},\dots,\mathsf{Ev}_{T})$: optionally produces a compact linkage proof (e.g., via hash chains or Merkle trees) for a batch of evidence records.
The original work [kao2025constant] focuses on classical adversaries, defining integrity and linkage properties and providing a concrete instantiation where each $\mathsf{Ev}$ is a 288- or 320-byte record. In this paper we treat $\mathsf{Generate}$ and $\mathsf{Verify}$ as abstract interfaces that we instantiate with post-quantum primitives and analyze against quantum adversaries.
2.2 Quantum Adversaries and Post-Quantum Primitives
We model adversaries as quantum polynomial-time (QPT) algorithms with access to certain oracles and system APIs. We write $\mathcal{A}$ for such adversaries and assume that they may maintain quantum side information throughout the experiment.
Quantum random-oracle model.
We adopt the quantum random-oracle model (QROM) for hash functions $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$, following the treatment of Zhandry and subsequent work [zhandry2012qrom, bonehzhandry2013qro]. In the QROM, $H$ is modeled as a uniformly random function, and the adversary may query it in quantum superposition. We assume that $H$ is collision-resistant against quantum adversaries, i.e., no QPT adversary can find distinct $x,x^{\prime}$ with $H(x)=H(x^{\prime})$ with non-negligible probability.
Post-quantum signature schemes.
We consider a signature scheme $\Sigma=(\mathsf{KeyGen}^{\Sigma},\mathsf{Sign}^{\Sigma},\mathsf{Verify}^{\Sigma})$ with standard syntax:
  • $(\mathsf{sk}^{\Sigma},\mathsf{pk}^{\Sigma})\leftarrow\mathsf{KeyGen}^{\Sigma}(1^{\lambda})$;
  • $\sigma\leftarrow\mathsf{Sign}^{\Sigma}(\mathsf{sk}^{\Sigma},m)$;
  • $b\leftarrow\mathsf{Verify}^{\Sigma}(\mathsf{pk}^{\Sigma},m,\sigma)$.
We assume that $\Sigma$ is existentially unforgeable under chosen-message attacks (EUF-CMA) against quantum adversaries: no QPT adversary with quantum access to $H$ and classical access to the signing oracle can produce a valid signature on a new message with non-negligible probability. We write $\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\lambda)$ for the corresponding advantage.
Concretely, $\Sigma$ may correspond to a lattice-based or hash-based scheme such as those standardized or under consideration in post-quantum cryptography efforts (e.g., [nistpqc, unruh2017pqsignatures]), but in this paper we treat it as an abstract primitive satisfying the above property. Real-world deployments face non-trivial challenges around key sizes, performance, and migration, as surveyed in [tan2022pqsignsurvey].
Notation.
We write $\lambda$ for the security parameter, $\mathsf{negl}(\lambda)$ for a negligible function, and use $\approx$ to denote computational indistinguishability. All probabilities are taken over the randomness of the algorithms and the random oracle $H$ unless otherwise specified.
3 Security Notions Against Quantum Adversaries
We now formalize security notions for evidence structures in the presence of quantum adversaries. Intuitively, we wish to prevent an adversary from forging evidence items for events that never happened, from equivocating about the content of the log, and from reusing the same evidence item for multiple conflicting events.
3.1 System Model with Quantum Adversaries
We consider an environment where an evidence generator uses $\mathsf{Generate}$ to produce evidence items for events and stores them in an append-only audit log. An auditor uses $\mathsf{Verify}$ and (optionally) $\mathsf{Link}$ to check evidence for consistency with a claimed workflow.
A quantum adversary $\mathcal{A}$ observes the log and may compromise the underlying storage or network. Specifically, we allow $\mathcal{A}$ to:
  • monitor and copy all evidence items produced during the lifetime of the system;
  • reorder, delete, or inject evidence items in the stored log;
  • interact with the evidence generator through a public API that exposes $\mathsf{Generate}$ for chosen events;
  • obtain quantum access to the random oracle $H$;
  • in a harvest-now, forge-later scenario, delay its forgery attempt until a future time when quantum resources become available.
We do not model side-channel attacks on key storage or the internals of the evidence generator; instead, we assume that the signing key is maintained inside a trusted environment (e.g., a hardware security module or a trusted execution environment). Extending our model to include side channels is an interesting direction for future work (Section 8).
3.2 Game-Based Definitions
We now define security games capturing the three main properties we seek: Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding.
Q-Audit Integrity.
Intuitively, Q-Audit Integrity requires that an adversary cannot produce a valid evidence item for an event that was not output by an honest execution of $\mathsf{Generate}$.
Definition 1 (Q-Audit Integrity).
Let $\mathcal{E}=(\mathsf{Setup},\mathsf{KeyGen},\mathsf{Generate},\mathsf{Verify},\mathsf{Link})$ be an evidence structure. Define the experiment $\mathsf{Exp}^{\mathsf{qaudit}}_{\mathcal{A}}(1^{\lambda})$ as follows:
  1. 1.
    The challenger runs $\mathsf{Setup}(1^{\lambda})$ to obtain $\mathsf{pp}$ and $\mathsf{KeyGen}(\mathsf{pp})$ to obtain $(\mathsf{sk},\mathsf{pk})$.
  2. 2.
    The adversary $\mathcal{A}$ is given $(\mathsf{pp},\mathsf{pk})$ and quantum access to the random oracle $H$. It also has classical access to a $\mathsf{Generate}$ oracle that, on input $e$, returns $\mathsf{Ev}\leftarrow\mathsf{Generate}(\mathsf{pp},\mathsf{sk},e)$ and adds $(e,\mathsf{Ev})$ to a set $\mathcal{Q}$ of generated pairs.
  3. 3.
    Eventually $\mathcal{A}$ outputs a pair $(e^{\star},\mathsf{Ev}^{\star})$.
  4. 4.
    The experiment outputs $1$ (meaning $\mathcal{A}$ wins) if and only if $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e^{\star},\mathsf{Ev}^{\star})=1$ and $(e^{\star},\mathsf{Ev}^{\star})\notin\mathcal{Q}$.
We define the Q-Audit Integrity advantage of $\mathcal{A}$ as $\mathsf{Adv}^{\mathsf{qaudit}}_{\mathcal{E}}(\mathcal{A},\lambda)=\Pr[\mathsf{Exp}^{\mathsf{qaudit}}_{\mathcal{A}}(1^{\lambda})=1]$. We say that $\mathcal{E}$ satisfies Q-Audit Integrity if for all QPT adversaries $\mathcal{A}$, $\mathsf{Adv}^{\mathsf{qaudit}}_{\mathcal{E}}(\mathcal{A},\lambda)$ is negligible in $\lambda$.
Q-Non-Equivocation.
Non-equivocation informally means that the adversary cannot present two mutually inconsistent views of the evidence log that both appear valid to honest auditors. There are several ways to formalize this; we adopt a simplified batch-based game.
Definition 2 (Q-Non-Equivocation).
Fix an ordering of events and let $\mathcal{L}$ denote a log, i.e., a sequence of pairs $(e_{i},\mathsf{Ev}_{i})$. We assume that the auditor interprets the $i$-th position of a log as the $i$-th logical step in the workflow, so conflicting entries at the same index correspond to equivocation about that step. Consider the experiment $\mathsf{Exp}^{\mathsf{qne}}_{\mathcal{A}}(1^{\lambda})$:
  1. 1.
    The challenger runs $\mathsf{Setup},\mathsf{KeyGen}$ as before and exposes $\mathsf{pp},\mathsf{pk}$ to $\mathcal{A}$.
  2. 2.
    $\mathcal{A}$ has oracle access to $\mathsf{Generate}$ and $H$, and may maintain an internal view of a log $\mathcal{L}$ obtained by querying $\mathsf{Generate}$ on events of its choice.
  3. 3.
    Eventually $\mathcal{A}$ outputs two logs $\mathcal{L}^{(0)}=\{(e^{(0)}_{i},\mathsf{Ev}^{(0)}_{i})\}_{i}$ and $\mathcal{L}^{(1)}=\{(e^{(1)}_{i},\mathsf{Ev}^{(1)}_{i})\}_{i}$.
  4. 4.
    The experiment checks:
    • for both $b\in\{0,1\}$, all pairs in $\mathcal{L}^{(b)}$ verify under $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},\cdot,\cdot)$; and
    • there exists at least one index $i$ such that $(e^{(0)}_{i},\mathsf{Ev}^{(0)}_{i})\neq(e^{(1)}_{i},\mathsf{Ev}^{(1)}_{i})$ but both are presented at the same log position $i$ in their respective views.
  5. 5.
    The experiment outputs $1$ if both bullets hold.
We define the Q-Non-Equivocation advantage $\mathsf{Adv}^{\mathsf{qne}}_{\mathcal{E}}(\mathcal{A},\lambda)$ accordingly, and say that $\mathcal{E}$ satisfies Q-Non-Equivocation if this advantage is negligible for all QPT adversaries.
The second bullet captures the notion that the two logs represent conflicting stories about the same workflow (e.g., a particular event is associated with different evidence items), while both appear individually valid.
Q-Binding.
Finally, we capture the requirement that an individual evidence item cannot be reused to attest to two distinct events.
Definition 3 (Q-Binding).
Define the experiment $\mathsf{Exp}^{\mathsf{qbind}}_{\mathcal{A}}(1^{\lambda})$:
  1. 1.
    The challenger runs $\mathsf{Setup},\mathsf{KeyGen}$ and gives $(\mathsf{pp},\mathsf{pk})$ and oracle access to $H$ to $\mathcal{A}$.
  2. 2.
    $\mathcal{A}$ may query a $\mathsf{Generate}$ oracle as before and obtains pairs $(e,\mathsf{Ev})$.
  3. 3.
    Eventually $\mathcal{A}$ outputs a triple $(e_{0},e_{1},\mathsf{Ev}^{\star})$ with $e_{0}\neq e_{1}$.
  4. 4.
    The experiment outputs $1$ if $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e_{0},\mathsf{Ev}^{\star})=\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e_{1},\mathsf{Ev}^{\star})=1$.
We define $\mathsf{Adv}^{\mathsf{qbind}}_{\mathcal{E}}(\mathcal{A},\lambda)$ as the success probability and say that $\mathcal{E}$ satisfies Q-Binding if this advantage is negligible for all QPT adversaries.
In the next section we show that a natural hash-and-sign instantiation using a PQ signature scheme achieves these properties under standard assumptions.
3.3 Relation to Classical Notions
In the special case where adversaries are classical and have only classical access to the hash oracle, our definitions reduce to classical notions of integrity and non-equivocation for secure logging and transparency systems [securelogging, ct]. Our experiments focus on the inability to create new valid evidence, to maintain two conflicting yet valid views of the log, and to rebind an evidence item. These correspond to the informal goals discussed in earlier work on constant-size evidence structures in a classical setting (e.g., [kao2025constant]) but here explicitly account for quantum-accessible oracles and quantum EUF-CMA adversaries.
4 Post-Quantum Instantiations of Evidence Structures
We now describe a concrete instantiation of $\mathsf{Generate}$ and $\mathsf{Verify}$ using a hash-and-sign paradigm in the quantum random-oracle model, and analyze its security with respect to the notions introduced above.
4.1 Hash-and-Sign Construction in the QROM
Let $\Sigma=(\mathsf{KeyGen}^{\Sigma},\mathsf{Sign}^{\Sigma},\mathsf{Verify}^{\Sigma})$ be a PQ signature scheme as in Section 2.2, and let $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$ be a hash function modeled as a random oracle in the QROM.
We assume that for each event $e$ we can deterministically compute a canonical encoding $\mathsf{Fields}(e)\in\{0,1\}^{*}$ that concatenates all fields to be included in the evidence record (e.g., code hash, input digest, platform measurement, timestamps). The precise layout of these fields follows the constant-size design from [kao2025constant]; we abstract it here as a bitstring.
Given this setup, we define:
  • $\mathsf{Setup}$ runs the underlying setup for $\Sigma$ and fixes $H$;
  • $\mathsf{KeyGen}$ runs $\mathsf{KeyGen}^{\Sigma}$ to produce $(\mathsf{sk}^{\Sigma},\mathsf{pk}^{\Sigma})$ and sets $(\mathsf{sk},\mathsf{pk})=(\mathsf{sk}^{\Sigma},\mathsf{pk}^{\Sigma})$;
  • $\mathsf{Generate}(\mathsf{pp},\mathsf{sk},e)$:
    1. 1.
      Compute $x\leftarrow\mathsf{Fields}(e)$.
    2. 2.
      Compute $m\leftarrow H(x)$.
    3. 3.
      Compute $\sigma\leftarrow\mathsf{Sign}^{\Sigma}(\mathsf{sk}^{\Sigma},m)$.
    4. 4.
      Output evidence $\mathsf{Ev}=(x,\sigma)$, padded or encoded as a constant-size record.
  • $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e,\mathsf{Ev})$:
    1. 1.
      Parse $\mathsf{Ev}$ as $(x,\sigma)$ and check that $x$ is a well-formed encoding of $\mathsf{Fields}(e)$. If not, return 0.
    2. 2.
      Compute $m\leftarrow H(x)$.
    3. 3.
      Return $\mathsf{Verify}^{\Sigma}(\mathsf{pk}^{\Sigma},m,\sigma)$.
In practice, $x$ may be compressed to a fixed-length representation by carefully designing the field layout; for our security analysis we only require that the mapping $e\mapsto x=\mathsf{Fields}(e)$ is deterministic and collision-resistant, in the sense that it is infeasible for a QPT adversary to find $e\neq e^{\prime}$ such that $\mathsf{Fields}(e)=\mathsf{Fields}(e^{\prime})$, where $e$ is understood as the full structured event description (including identifiers, digests, and metadata).
4.2 Post-Quantum Signature Choices
Our analysis is agnostic to the concrete choice of $\Sigma$ as long as it satisfies quantum EUF-CMA security. In practice, the choice of signature scheme significantly affects the size of $\sigma$ and the performance of $\mathsf{Sign}$ and $\mathsf{Verify}$. For example, lattice-based signatures typically offer relatively short signatures and fast verification at the cost of larger public keys, while hash-based schemes offer conservative security at the cost of larger signatures or statefulness.
From the perspective of constant-size evidence records, the signature size directly contributes to the overall record size. If the base layout for $x$ is, say, $160$ bytes (including digests and metadata) and the signature is $128$ bytes, the resulting record size is $288$ bytes (plus any padding or alignment required), which is in the same range as prior designs [kao2025constant]. The throughput impact can be estimated by benchmarking the chosen $\Sigma$ on representative hardware; our case study in Section 6 sketches such an analysis.
4.3 Security Theorems
We now state our main security theorem for the hash-and-sign instantiation. Proofs follow standard reduction techniques from signature and QROM analyses; we provide sketches here and defer full details to an extended version.
Theorem 1 (Q-Audit Integrity and Q-Binding).
Assume that $H$ is collision-resistant against quantum adversaries in the QROM and that $\Sigma$ is EUF-CMA secure against quantum adversaries. Then the hash-and-sign evidence structure described above satisfies Q-Audit Integrity and Q-Binding in the sense of the experiments $\mathsf{Exp}^{\mathsf{qaudit}}$ and $\mathsf{Exp}^{\mathsf{qbind}}$.
Proof sketch.
We outline the argument for Q-Audit Integrity; the Q-Binding property is similar.
Suppose there exists a QPT adversary $\mathcal{A}$ that wins $\mathsf{Exp}^{\mathsf{qaudit}}_{\mathcal{A}}(1^{\lambda})$ with non-negligible probability $\epsilon(\lambda)$. We construct a QPT adversary $\mathcal{B}$ that breaks either the EUF-CMA security of $\Sigma$ or the collision resistance of $H$.
Adversary $\mathcal{B}$ is given a public key $\mathsf{pk}^{\Sigma}$ and oracle access to $\mathsf{Sign}^{\Sigma}$ and $H$ in the QROM. It sets $\mathsf{pk}=\mathsf{pk}^{\Sigma}$ and simulates the $\mathsf{Generate}$ oracle for $\mathcal{A}$ by: on input $e$, computing $x=\mathsf{Fields}(e)$, $m=H(x)$, and obtaining $\sigma\leftarrow\mathsf{Sign}^{\Sigma}(m)$ from its signing oracle, then returning $\mathsf{Ev}=(x,\sigma)$. It maintains the set $\mathcal{Q}$ of generated pairs as in the experiment. Note that $\mathcal{B}$ can answer all of $\mathcal{A}$’s hash queries by forwarding them to the QROM oracle.
Eventually, $\mathcal{A}$ outputs $(e^{\star},\mathsf{Ev}^{\star})$ with $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ such that $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e^{\star},\mathsf{Ev}^{\star})=1$ and $(e^{\star},\mathsf{Ev}^{\star})\notin\mathcal{Q}$. There are two cases:
  1. 1.
    If $x^{\star}$ equals $\mathsf{Fields}(e)$ for some event $e$ for which $\mathcal{B}$ already obtained a signature via the signing oracle, then $(e^{\star},\mathsf{Ev}^{\star})\in\mathcal{Q}$, contradicting the success condition of $\mathcal{A}$. Thus, this case can only occur with negligible probability due to simulation inconsistencies, which can be bounded using standard QROM techniques.
  2. 2.
    Otherwise, either $x^{\star}$ is a new input to the signing oracle or $\sigma^{\star}$ is a new signature on a previously queried $m^{\star}=H(x^{\star})$. In the former case, $(x^{\star},\sigma^{\star})$ constitutes a valid forgery against the EUF-CMA security of $\Sigma$; $\mathcal{B}$ outputs $(m^{\star},\sigma^{\star})$ as its forgery. In the latter case, if $\mathcal{B}$ did not query $H$ on $x^{\star}$, we can use standard techniques to extract a collision in $H$ from $\mathcal{A}$’s behavior in the QROM, contradicting collision resistance.
More formally, we construct $\mathcal{B}$ as a sequence of hybrids, gradually replacing oracle responses with simulated ones and bounding the statistical distance between experiments. The final hybrid yields either a signature forgery or a hash collision with probability non-negligibly related to $\epsilon(\lambda)$. Hence, if $\mathsf{Adv}^{\mathsf{qaudit}}_{\mathcal{E}}(\mathcal{A},\lambda)$ were non-negligible, one of $\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\lambda)$ or the collision-resistance advantage for $H$ would be non-negligible, contradicting our assumptions.
A fully rigorous proof in the QROM requires a careful treatment of quantum oracle simulation and programming; we omit these standard technical details here and refer to, e.g., [zhandry2012qrom, bonehzhandry2013qro, unruh2017pqsignatures] for the underlying techniques. ∎
A similar reduction shows Q-Binding: if an adversary can produce a single evidence item $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ that verifies as valid for two distinct events $e_{0},e_{1}$, then either $\mathsf{Fields}(e_{0})=\mathsf{Fields}(e_{1})$ (which we assume cannot happen for distinct events) or $\mathsf{Fields}(e_{0})\neq\mathsf{Fields}(e_{1})$, in which case the adversary has produced two different events with the same $x^{\star}$, contradicting the collision-resistance of the mapping $e\mapsto\mathsf{Fields}(e)$ (or of $H$ if $x$ is compressed further). This corresponds to the usual binding intuition: a single authenticator cannot be successfully opened as two different well-formed event descriptions except with negligible probability.
Non-equivocation can be shown under additional assumptions about the linkage mechanism (e.g., hash chains or Merkle trees) that prevent reordering and omission of evidence items, following techniques from transparency logs [ct]. We conjecture that Q-Non-Equivocation holds under standard assumptions when the linkage mechanism is instantiated with a collision-resistant hash function in the QROM; a full formalization and proof is left as future work.
5 Quantum-Safe Migration of Deployed Evidence Logs
We now turn from the design of new evidence structures to the question of how to migrate already-deployed evidence logs towards quantum safety. We assume the existence of a system that has been using a classical signature scheme to protect constant-size evidence records and that must remain auditable for a long period (e.g., 10–30 years) in the face of emerging quantum capabilities.
5.1 Threat Scenarios: Harvest-Now, Forge-Later
A particularly relevant threat scenario for audit logs is the harvest-now, forge-later model: an adversary records all publicly visible evidence items today and stores them. Once a sufficiently powerful quantum computer becomes available, the adversary attempts to:
  • break the classical signature scheme (e.g., via Shor’s algorithm) to produce valid signatures on arbitrary messages;
  • or use the ability to forge signatures to create false evidence records or to alter existing ones in the stored log;
  • or present two conflicting views of the evidence log to different auditors.
In this scenario, the adversary does not need to compromise the system in real time: passive eavesdropping and later cryptanalysis suffice. This distinguishes audit migration from other settings (e.g., TLS) where the focus is on protecting present and future sessions rather than long-lived, highly replicable logs. In contrast to PQ-TLS migration, which focuses on protecting future sessions [alnahawi2024pqtls], audit trails must preserve the evidentiary value of past executions.
If no migration is performed, any evidence record that relies solely on classical signatures may lose its probative value once the relevant scheme is broken: a skeptical auditor could no longer trust that a signature was produced at the claimed time or by the claimed key holder. Our goal is therefore to design migration strategies that augment or wrap existing evidence with post-quantum protections in a way that preserves their evidentiary meaning.
5.2 Migration Patterns
We discuss three migration patterns, each suitable for different operational and regulatory constraints.
Pattern 1: Hybrid signatures for new evidence.
The simplest migration pattern is to augment all new evidence records with a post-quantum signature in addition to the existing classical one. Concretely, for each new event $e$ we generate:
$\mathsf{Ev}^{\mathrm{hyb}}=(x,\sigma^{\mathrm{class}},\sigma^{\mathrm{pq}}),$
where $x=\mathsf{Fields}(e)$, $\sigma^{\mathrm{class}}$ is a signature under the existing classical scheme (e.g., ECDSA), and $\sigma^{\mathrm{pq}}$ is a signature under a PQ scheme $\Sigma$. Verification checks both signatures.
Security-wise, as long as either signature scheme remains secure, an adversary cannot forge a full hybrid evidence item. In particular, once the classical scheme is broken by quantum computers, the PQ signature continues to provide integrity and binding. The main cost is increased record size and verification time due to the additional signature; in systems where records are already small (e.g., a few hundred bytes), the overhead may be acceptable.
Pattern 2: Re-signing legacy evidence.
Hybrid signatures do not protect evidence generated before a PQ scheme is deployed. For legacy records, one option is to re-sign them with a PQ key inside a trusted environment. Let $\{\mathsf{Ev}_{i}\}$ denote legacy evidence records containing classical signatures. The system operator can:
  1. 1.
    deploy a trusted re-signing service (e.g., in a hardware security module or TEE) with a PQ signing key $\mathsf{sk}^{\mathrm{pq}}$;
  2. 2.
    for each legacy record $\mathsf{Ev}_{i}$, compute a digest $d_{i}=H^{\prime}(\mathsf{Ev}_{i})$ under a quantum-resistant hash $H^{\prime}$;
  3. 3.
    compute a PQ signature $\tau_{i}=\mathsf{Sign}^{\Sigma}(\mathsf{sk}^{\mathrm{pq}},d_{i})$;
  4. 4.
    store the pair $(\mathsf{Ev}_{i},\tau_{i})$ as the migrated version of the record.
Auditors can then verify both the original classical evidence and its PQ wrapper: $\mathsf{Verify}^{\Sigma}(\mathsf{pk}^{\mathrm{pq}},H^{\prime}(\mathsf{Ev}_{i}),\tau_{i})$. The trust assumption is that the re-signing process is performed honestly and that the original log was unmodified at the time of re-signing. This process effectively extends the lifetime of legacy evidence by anchoring it to a PQ signature scheme in a single step.
Pattern 3: Merkle-root anchoring for legacy batches.
When the number of legacy evidence items is very large, re-signing each item individually may be computationally expensive. An alternative is to anchor batches of legacy records via Merkle trees. For a batch $B=\{\mathsf{Ev}_{1},\dots,\mathsf{Ev}_{n}\}$, the system:
  1. 1.
    builds a Merkle tree $T_{B}$ where the leaves are $h_{i}=H^{\prime}(\mathsf{Ev}_{i})$;
  2. 2.
    computes the root $r_{B}$ of $T_{B}$;
  3. 3.
    generates a PQ signature $\tau_{B}=\mathsf{Sign}^{\Sigma}(\mathsf{sk}^{\mathrm{pq}},r_{B})$;
  4. 4.
    stores $\tau_{B}$ as the PQ anchor for batch $B$.
To verify a legacy record $\mathsf{Ev}_{i}$ in $B$, an auditor obtains a Merkle proof $\pi_{i}$ that $\mathsf{Ev}_{i}$ is included in $T_{B}$, verifies the proof against $r_{B}$, and verifies the PQ signature on $r_{B}$. This pattern amortizes the cost of PQ signatures over many records; the trade-off is the need to manage Merkle proofs and batch boundaries.
5.3 Complexity and Storage Analysis
We now sketch the rough complexity of these patterns in terms of the number of legacy records $N$, the size of a PQ signature $|\sigma^{\mathrm{pq}}|$, and the batch size $b$ for Merkle anchoring.
Hybrid signatures.
For new evidence only, the cost is one additional PQ signature per record. If we generate $N_{\mathrm{new}}$ evidence items over the remaining lifetime of the system, the total additional signing cost is $N_{\mathrm{new}}$ PQ signatures; storage overhead is $N_{\mathrm{new}}\cdot|\sigma^{\mathrm{pq}}|$ bytes.
Re-signing legacy evidence.
Re-signing all $N$ legacy records individually requires $N$ PQ signatures and adds $N\cdot|\tau_{i}|$ bytes of storage, where $|\tau_{i}|$ is the size of the PQ signature (and possibly some metadata). The process is embarrassingly parallel and can be spread over time, but for very large $N$ this may still be expensive.
Merkle-root anchoring.
If we partition $N$ records into batches of size $b$, we obtain $\lceil N/b\rceil$ roots and therefore need only $\lceil N/b\rceil$ PQ signatures. The storage overhead for signatures is thus $\lceil N/b\rceil\cdot|\tau_{B}|$. Each record also requires storing or reconstructing a Merkle proof of size $O(\log b)$, e.g., a few dozen hashes for typical batch sizes. The one-time computation cost is dominated by hashing all records ($N$ hashes) and signing $\lceil N/b\rceil$ roots.
Table 1 summarizes the trade-offs qualitatively.
Pattern PQ signatures Per-record overhead Trust assumptions
Hybrid (new only) $N_{\mathrm{new}}$ $|\sigma^{\mathrm{pq}}|$ Same as existing system
Re-sign legacy NN $|\tau_{i}|$ Trusted re-signing process
Merkle-root anchoring $\lceil N/b\rceil$ $O(\log b)$ hashes Trusted tree construction
Table 1: Qualitative comparison of migration patterns for $N$ legacy evidence records.
In practice, combinations of these patterns may be used: for example, Merkle-root anchoring for very old archival logs and full re-signing for recent, higher-value records. Recent system designs that combine PQC and blockchain for audit-grade file transfer [sola2025qrfile] follow a complementary pattern, anchoring events in an immutable ledger rather than constant-size evidence records.
6 Case Study: A Constant-Size Evidence System in Regulated AI
To illustrate how the above concepts translate into practice, we sketch a case study based on a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc. The platform targets clinical and pharmaceutical workflows where each AI job (e.g., a risk prediction or trial cohort selection) is executed inside a hardened containerized environment and produces a compact evidence record.
6.1 Deployment Overview
In a typical deployment, the platform orchestrates AI jobs across a cluster of container-based trusted execution environments (TEEs). For each job, the runtime collects:
  • a hash of the container image (code and model);
  • digests of input datasets and configuration;
  • a platform measurement from the TEE (e.g., a measurement of the guest VM);
  • wall-clock timestamps and policy identifiers.
These fields are combined into a constant-size record $x=\mathsf{Fields}(e)$ and signed to produce an evidence item $\mathsf{Ev}=(x,\sigma)$, following the abstraction in Section 2.1. Evidence items are stored in an append-only log backed by a database; auditors can later retrieve evidence for any job and verify it against the public key of the platform.
In current deployments, the signature scheme is classical (e.g., ECDSA) while the rest of the design (constant-size layout, hash-based linkage) is agnostic to the choice of signature. The audit log may grow to $10^{6}$$10^{8}$ evidence records over several years, depending on workload volume.
6.2 Applying Migration Patterns
We now consider how the migration patterns from Section 5.2 apply in this setting.
Hybrid signatures for new evidence.
Once a PQ signature scheme is available in the platform’s cryptographic module, new jobs can produce hybrid evidence items $(x,\sigma^{\mathrm{class}},\sigma^{\mathrm{pq}})$ without changing the upstream workflow. The primary engineering task is to extend the evidence generator and verifier to handle an additional signature field. Because evidence records are already compact (e.g., $\approx 288$ bytes), adding a PQ signature of similar size roughly doubles the record size but keeps it under 1 KB, which is acceptable for many log storage systems.
Re-signing recent legacy evidence.
For recent evidence, especially those corresponding to higher-risk workflows (e.g., clinical trial analyses), the platform can re-sign individual records using a PQ signature key inside a TEE. For concreteness, we assume a PQ signature throughput of roughly 5,0005{,}000 signatures per second on commodity CPUs, in line with reported benchmarks for lattice-based signatures in TLS and PKI settings [raavi2025pqdsa]. Given $N_{\mathrm{recent}}$ such records, re-signing $10^{7}$ records would take on the order of 2,000 seconds ($\approx 34$ minutes) on a single core, or significantly less when parallelized. These back-of-the-envelope estimates suggest that re-signing at this scale is operationally feasible.
Merkle-root anchoring for archival logs.
For older archival logs, where individual records are rarely accessed but must remain provably intact, Merkle-root anchoring can reduce migration cost. For example, with batch size $b=2^{12}=4{,}096$, a corpus of $N=10^{8}$ records yields $\lceil N/b\rceil\approx 24{,}415$ roots. At a PQ signing throughput of 5,0005{,}000 signatures per second, computing all anchors would take under 5 seconds of CPU time; the dominant cost lies in computing the $N$ hash values for the leaves, which is linear in the number of records and can be parallelized. Merkle proofs of depth $\log_{2}b=12$ hashes add modest per-record overhead.
6.3 Lessons for Regulated AI Systems
This case study highlights several lessons:
  • Designing evidence records to be cryptographically modular—i.e., not hard-wiring a specific signature scheme into their layout—greatly simplifies PQ migration.
  • For realistic log volumes in regulated AI workloads, the computational cost of PQ migration is moderate compared to typical batch processing tasks.
  • Hybrid and batch anchoring strategies enable operators to prioritize which portions of the log receive the strongest PQ protection, aligning cryptographic investment with regulatory and business risk.
While the concrete numbers above are illustrative rather than definitive, they suggest that the migration patterns we propose can be implemented without fundamental scalability obstacles.
7 Related Work
Secure logging and transparency.
Classical secure logging and forward-secure audit log designs (e.g., [securelogging, forwardsecurelogging, haberstornetta1991, merkle1989certified]) focus on protecting logs against classical adversaries, often using hash chains and forward-secure keys. Transparency logs such as Certificate Transparency and Revocation Transparency [ct, transparency] add public append-only structures and gossip protocols to detect misbehavior. Our work builds on these foundations but explicitly targets quantum adversaries and constant-size evidence structures for regulated AI settings.
Post-quantum cryptography and logging.
Post-quantum cryptography has been extensively studied and standardized, e.g., [bernstein2009pqcbook, nistpqc, bonehshoup]. More recently, several works have explored quantum-resilient logging and auditing in domain-specific settings, such as distributed logs [somarapu2022qrlogs], cloud multimedia auditing [khan2025pqblockchainaudit], and telecom logs for 5G/6G [yusuf2025qrlogintegrity]. Survey data on PQC support in mainstream cryptographic libraries [ahmed2025pqclibs] indicates that hybrid and wrapper-based migration, as in our patterns, aligns well with how PQ algorithms are currently being integrated into software stacks. Our work is complementary: we focus on constant-size evidence structures and provide a unified treatment of quantum-adversary notions and migration patterns for regulated AI audit trails.
8 Discussion and Open Problems
Our work raises several broader questions about designing quantum-safe audit trails.
Choice of PQ primitives.
Different PQ signature schemes offer different trade-offs among key size, signature size, signing and verification speed, and implementation maturity. Our analysis treats $\Sigma$ as an abstract EUF-CMA-secure primitive, but deploying such schemes in evidence systems requires careful engineering and benchmarking. In particular, regulators may favor conservative, hash-based schemes despite larger signatures, while performance-sensitive environments may prefer lattice-based schemes.
Policy and retention considerations.
Cryptographic migration is only one part of a broader policy question: for how long must evidence be retained, at what integrity level, and under which key rotation schedule? For example, a regulator may consider it acceptable to rely on classical signatures for low-risk workflows with short retention periods, while demanding PQ protection for long-lived, high-impact logs. Formalizing such policies and aligning them with technical migration patterns is an interesting interdisciplinary challenge.
Beyond integrity: confidentiality and privacy.
In this paper we focus on integrity, non-repudiation, and binding. Many regulated AI settings also require protecting the confidentiality of certain log fields and respecting privacy constraints. Extending evidence structures to include post-quantum encryption (e.g., via KEM+AEAD constructions) and differential privacy mechanisms, while maintaining verifiability, is a promising direction.
Stronger quantum security models.
Our security model accounts for quantum access to hash oracles and EUF-CMA adversaries against signatures. In more advanced settings, adversaries may have coherent access to additional system APIs, or evidence structures may be embedded in protocols that themselves involve quantum communication. Developing stronger compositional security frameworks for such scenarios remains an open problem.
Side channels and implementation security.
We do not consider side-channel attacks or implementation bugs that leak signing keys or allow evidence tampering. While these issues are orthogonal to PQ migration, they are critical in practice. Combining post-quantum cryptography with formally verified implementations and side-channel-resistant TEEs would further strengthen audit trails.
9 Conclusion
We have studied evidence structures for regulated AI audit trails in the presence of quantum adversaries. Building on an abstraction of constant-size evidence records, we introduced game-based notions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, and analyzed a hash-and-sign instantiation in the quantum random-oracle model using post-quantum signature schemes. We then proposed migration patterns for deployed evidence logs—hybrid signatures, re-signing of legacy records, and Merkle-root anchoring—and examined their trade-offs.
Our case study, based on an industrial constant-size evidence platform for regulated AI at Codebat Technologies Inc., suggests that quantum-safe audit trails are practically achievable with moderate computational and storage overhead. We hope that our definitions and patterns provide a useful foundation for both researchers and practitioners designing quantum-resilient compliance systems.
Future work includes extending our analysis to richer linkage mechanisms, integrating confidentiality and privacy requirements, and collaborating with regulators to align quantum-safe audit designs with evolving standards.

### v2

@@
-Quantum-Adversary-Resilient Evidence Structures and Migration Strategies for Regulated AI Audit Trails
+Post-Quantum-Resilient Audit Evidence for Long-Lived Regulated Systems: Security Models, Migration Patterns, and Case Study
 Abstract
@@
 We recall an abstraction of constant-size evidence structures and introduce game-based definitions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, capturing the inability of a quantum adversary to forge, equivocate, or rebind evidence items.
-We then analyze a hash-and-sign instantiation in the quantum random-oracle model (QROM), assuming an existentially unforgeable PQ signature scheme against quantum adversaries, and show that the resulting evidence structure satisfies these notions under standard assumptions.
+We then analyze a hash-and-sign instantiation in the quantum random-oracle model (QROM), assuming an existentially unforgeable PQ signature scheme against quantum adversaries, and prove via tight reductions that the resulting evidence structure satisfies all three notions under standard assumptions (collision-resistant hashing and quantum EUF-CMA signatures).
 Building on this, we present three migration patterns for existing evidence logs—hybrid signatures, re-signing of legacy evidence, and Merkle-root anchoring—and analyze their security, storage, and computational trade-offs.
-A case study based on an industrial constant-size evidence platform for regulated AI at Codebat Technologies Inc.
-suggests that quantum-safe audit trails are achievable with moderate overhead and that systematic migration can significantly extend the evidentiary lifetime of existing deployments.
+A case study based on an industrial constant-size evidence platform for regulated AI at Codebat Technologies Inc., implemented in Rust and benchmarked on commodity hardware, suggests that quantum-safe audit trails are achievable with moderate overhead and that systematic migration can significantly extend the evidentiary lifetime of existing deployments.
 1 Introduction
@@
 In domains such as clinical AI, pharmaceutical development, and high-stakes financial analytics, regulators require that organizations be able to explain and justify how a particular result was produced: which code and model were used, which data were processed, under which policies, on which platform, and at what time.
-Recent system designs address these needs by representing each execution as a constant-size cryptographic evidence item that compactly captures the essential provenance and integrity information and can be stored and verified at scale (e.g., [kao2025constant, securelogging, transparency]), and recent work has begun to study quantum-resilient designs for distributed log storage itself [somarapu2022qrlogs].
+Recent system designs address these needs by representing each execution as a constant-size cryptographic evidence item that compactly captures the essential provenance and integrity information and can be stored and verified at scale (e.g., [7, 14, 9]), and recent work has begun to study quantum-resilient designs for distributed log storage itself [16].
 Today, many such systems instantiate their evidence structures with classical primitives such as ECDSA or RSA signatures and SHA-2 family hash functions.
@@
 In this work we address these gaps.
-We build on an abstraction of constant-size evidence structures introduced in prior work [kao2025constant], which models a system where each event in a workflow is associated with a fixed-length evidence record produced by a Generate algorithm and consumed by a Verify algorithm.
+We build on an abstraction of constant-size evidence structures introduced in prior work [7], which models a system where each event in a workflow is associated with a fixed-length evidence record produced by a Generate algorithm and consumed by a Verify algorithm.
 Our focus here is on extending this abstraction to a quantum-adversary setting, providing both formal definitions and practically oriented migration strategies.
@@
   2. 2.
-    Post-quantum instantiations in the QROM.
-We analyze a hash-and-sign instantiation of constant-size evidence structures in the quantum random-oracle model, assuming an underlying post-quantum signature scheme that is existentially unforgeable under chosen-message attacks (EUF-CMA) against quantum adversaries.
-We prove that under standard assumptions the instantiated scheme satisfies our quantum-adversary security notions.
+    Post-quantum instantiations with complete proofs in the QROM.
+We analyze a hash-and-sign instantiation of constant-size evidence structures in the quantum random-oracle model and provide three complete reduction proofs with explicit advantage bounds, showing that Q-Audit Integrity, Q-Binding, and Q-Non-Equivocation all hold under collision-resistant hashing and quantum EUF-CMA signatures.
   3. 3.
@@
   4. 4.
-    Industrial case study.
-We outline a case study using a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc.
-We discuss how the proposed migration patterns map to realistic deployment constraints and estimate the cost of securing large evidence corpora.
+    Industrial case study with PQ benchmarks.
+We outline a case study using a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc., including concrete performance comparisons across post-quantum signature schemes (ML-DSA, SLH-DSA) to quantify migration overhead.
+  5. 5.
+    Systematic comparison. We provide a feature-level comparison with existing post-quantum logging and audit systems, identifying the gaps that our work addresses.
 The rest of the paper is organized as follows.
 Section 2 recalls the abstraction of constant-size evidence structures and reviews quantum-adversary and PQ primitive models.
+Section 3 describes the deployment context and constant-size evidence layout in our industrial platform.
 Section 4 introduces our quantum-adversary security notions.
@@
 Relation to prior work.
-Compared to classical secure logging and transparency systems (e.g., [securelogging, forwardsecurelogging, haberstornetta1991, merkle1989certified, ct, transparency]), our work treats constant-size evidence structures as the primary cryptographic object and explicitly analyzes their security against quantum adversaries in the QROM.
-Existing PQC work on logging and auditing (e.g., [somarapu2022qrlogs, khan2025pqblockchainaudit, yusuf2025qrlogintegrity]) typically either focuses on blockchain-based storage or on specific application domains.
+Compared to classical secure logging and transparency systems [14, 3, 6, 11, 10, 9], our work treats constant-size evidence structures as the primary cryptographic object and explicitly analyzes their security against quantum adversaries in the QROM.
+Existing PQC work on logging and auditing (e.g., [16, 8, 19]) typically either focuses on blockchain-based storage or on specific application domains.
 In contrast, we provide (i) generic quantum-adversary security notions for evidence structures, and (ii) migration patterns that are designed to retrofit already-deployed constant-size evidence logs in regulated AI systems.
@@
 2.1 Constant-Size Evidence Structures
-We briefly recall the abstraction of constant-size evidence structures introduced in [kao2025constant], which we treat here as a black-box building block.
+We briefly recall the abstraction of constant-size evidence structures introduced in [7], which we treat here as a black-box building block.
 Consider a workflow that generates a sequence of events $e_{1},e_{2},\dots,e_{T}$, where each event $e_{i}$ corresponds to a concrete execution step such as a model inference, a batch data transformation, or an approval action.
@@
   • $\pi\leftarrow\mathsf{Link}(\mathsf{pp},\mathsf{Ev}_{1},\dots,\mathsf{Ev}_{T})$: optionally produces a compact linkage proof (e.g., via hash chains or Merkle trees) for a batch of evidence records.
-The original work [kao2025constant] focuses on classical adversaries, defining integrity and linkage properties and providing a concrete instantiation where each $\mathsf{Ev}$ is a 288- or 320-byte record.
+The original work [7] focuses on classical adversaries, defining integrity and linkage properties and providing a concrete instantiation where each $\mathsf{Ev}$ is a 288- or 320-byte record.
 In this paper we treat $\mathsf{Generate}$ and $\mathsf{Verify}$ as abstract interfaces that we instantiate with post-quantum primitives and analyze against quantum adversaries.
@@
 Quantum random-oracle model.
-We adopt the quantum random-oracle model (QROM) for hash functions $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$, following the treatment of Zhandry and subsequent work [zhandry2012qrom, bonehzhandry2013qro].
+We adopt the quantum random-oracle model (QROM) for hash functions $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$, following the treatment of Zhandry and subsequent work [20, 5].
 In the QROM, $H$ is modeled as a uniformly random function, and the adversary may query it in quantum superposition.
@@
 Post-quantum signature schemes.
-We consider a signature scheme $\Sigma=(\mathsf{KeyGen}^{\Sigma},\mathsf{Sign}^{\Sigma},\mathsf{Verify}^{\Sigma})$ with standard syntax:
+We consider a signature scheme $\Sigma$ with standard syntax:
   • $(\mathsf{sk}^{\Sigma},\mathsf{pk}^{\Sigma})\leftarrow\mathsf{KeyGen}^{\Sigma}(1^{\lambda})$;
@@
 We write $\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\lambda)$ for the corresponding advantage.
-Concretely, $\Sigma$ may correspond to a lattice-based or hash-based scheme such as those standardized or under consideration in post-quantum cryptography efforts (e.g., [nistpqc, unruh2017pqsignatures]), but in this paper we treat it as an abstract primitive satisfying the above property.
-Real-world deployments face non-trivial challenges around key sizes, performance, and migration, as surveyed in [tan2022pqsignsurvey].
+Concretely, $\Sigma$ may correspond to a lattice-based or hash-based scheme such as those standardized or under consideration in post-quantum cryptography efforts (e.g., [12, 18]), but in this paper we treat it as an abstract primitive satisfying the above property.
+Real-world deployments face non-trivial challenges around key sizes, performance, and migration, as surveyed in [17].
 Notation.
@@
 All probabilities are taken over the randomness of the algorithms and the random oracle $H$ unless otherwise specified.
+3 System Context: Constant-Size Evidence for Regulated AI
+Before introducing our quantum-adversary security model, we describe the deployment context and the constant-size evidence structure that motivates our formal treatment.
+This section draws on our prior work on constant-size evidence packs for regulated AI [7] and summarizes the key system design choices relevant to quantum-resilient migration.
+3.1 Deployment Scenario
+We consider regulated AI deployments in domains such as clinical decision support, pharmaceutical development, and financial analytics, where each AI workload execution must be auditable for compliance and forensic purposes.
+In a typical deployment, the platform orchestrates AI jobs across a cluster of container-based trusted execution environments (TEEs).
+For each job, the runtime collects attestation and provenance attributes and produces a constant-size evidence record that is appended to an audit log.
+Unlike short-lived session keys in TLS, audit evidence for regulated AI workloads is meant to support long-term accountability and forensic investigations.
+Hospitals, financial institutions, and pharmaceutical companies routinely retain audit logs for ten years or longer due to regulatory requirements.
+This creates an attractive “harvest-now, forge-later” target: an adversary can exfiltrate classical evidence logs today and exploit future quantum capabilities to retroactively forge or repudiate past AI executions.
+3.2 Evidence Record Layout
+Each AI job produces a constant-size evidence record capturing the key attestation and provenance attributes needed for regulated audit trails.
+Table 1 summarizes a representative layout used in our industrial deployment at Codebat Technologies Inc.
+Table 1: Representative constant-size evidence layout in our regulated AI deployment, following the abstraction from [7]. Each field is a $\lambda$-bit hash output (typically $\lambda=256$).
+Field Description Size (bits)
+$f_{0}$: context Workflow/job identifier commitment 256
+$f_{1}$: model Hash of model binary / parameters 256
+$f_{2}$: code Hash of container image / code snapshot 256
+$f_{3}$: input Hash of input dataset or configuration 256
+$f_{4}$: output Hash of output artifacts 256
+$f_{5}$: policy Policy state / IRB identifier commitment 256
+$f_{6}$: environment TEE measurement / platform attestation 256
+$f_{7}$: link Link to previous evidence or audit chain 256
+$\sigma$: signature Ed25519 or PQ signature 512–2048
+Total ($k=8$ fields + signature) 320–512 bytes
+The fixed-size layout ensures that each evidence record occupies the same storage footprint regardless of the complexity of the underlying AI job, enabling predictable storage costs and $O(1)$ per-record verification.
+The signature field currently uses a classical scheme (e.g., ECDSA or Ed25519) but is designed to accommodate post-quantum signatures as part of the migration strategies discussed in Section 6.
+3.3 Implementation and Baseline Performance
+We have implemented the constant-size evidence generator in Rust as part of an industrial platform for regulated AI deployments [7].
+The implementation uses a standard 256-bit hash function (SHA-256) and an Edwards-curve-based signature scheme (Ed25519).
+The number of fields $k$ is fixed at system setup time; in practice, $k\in[8,12]$ is sufficient for most clinical and pharmaceutical workflows.
+[…]
+Table 2: Evidence generation and verification throughput on commodity hardware.
+[…]
+Generation (single-threaded) $3.5\times 10^{4}$ 28.4 $k$ fixed, $\lambda=256$
+Generation (16 CPU cores) $2.8\times 10^{5}$ 5.7 near-linear scaling
+Verification (16 CPU cores) $2.5\times 10^{5}$ 6.1 batch of $10^{5}$ items
+Verification (GPU) $4.0\times 10^{5}$ 2.5 batch of $10^{6}$ items
+The fixed-size layout of evidence items leads to regular memory access patterns and uniform control flow, which are favorable for both vectorized CPU execution and GPU kernels.
+This $O(1)$ per-event cost is crucial for long-term retention, as regulated environments are expected to keep audit logs for 10–30 years while enforcing strict storage and verification latency budgets.
+A deployment handling $10^{3}$ AI jobs per day accumulates on the order of $10^{6}$$10^{7}$ evidence records over a decade, all of which must remain efficiently verifiable throughout their retention period.
 4 Security Notions Against Quantum Adversaries
@@
 Intuitively, we wish to prevent an adversary from forging evidence items for events that never happened, from equivocating about the content of the log, and from reusing the same evidence item for multiple conflicting events.
-3.1 System Model with Quantum Adversaries
-We consider an environment where an evidence generator uses $\mathsf{Generate}$ to produce evidence items for events and stores them in an append-only audit log.
+4.1 System and Threat Model
+We consider a regulated AI deployment where each execution of an AI workload produces a constant-size evidence record, as described in Section 3, and these records are appended to an audit log that is expected to be stored and queried over many years.
+The evidence generator is integrated with the AI platform’s scheduling and container orchestration layer, while one or more auditors rely on the evidence log to reconstruct and verify past executions during compliance checks and incident response.
+Formally, an evidence generator uses $\mathsf{Generate}$ to produce evidence items for events and stores them in an append-only audit log.
 An auditor uses $\mathsf{Verify}$ and (optionally) $\mathsf{Link}$ to check evidence for consistency with a claimed workflow.
@@
     The experiment checks:
-    • for both $b\in\{0,1\}$, all pairs in $\mathcal{L}^{(b)}$ verify under $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},\cdot,\cdot)$; and
-    • there exists at least one index $i$ such that $(e^{(0)}_{i},\mathsf{Ev}^{(0)}_{i})\neq(e^{(1)}_{i},\mathsf{Ev}^{(1)}_{i})$ but both are presented at the same log position $i$ in their respective views.
+    • for both $b\in\{0,1\}$, all pairs in $\mathcal{L}^{(b)}$ verify under $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},\cdot,\cdot)$;
+    • if $\mathsf{Link}$ is defined, $\mathsf{Link}(\mathsf{pp},\mathcal{L}^{(0)})=\mathsf{Link}(\mathsf{pp},\mathcal{L}^{(1)})$; and
+    • there exists at least one index $i$ such that $(e^{(0)}_{i},\mathsf{Ev}^{(0)}_{i})\neq(e^{(1)}_{i},\mathsf{Ev}^{(1)}_{i})$.
   5. 5.
-    The experiment outputs $1$ if both bullets hold.
+    The experiment outputs $1$ if all conditions hold.
 We define the Q-Non-Equivocation advantage $\mathsf{Adv}^{\mathsf{qne}}_{\mathcal{E}}(\mathcal{A},\lambda)$ accordingly, and say that $\mathcal{E}$ satisfies Q-Non-Equivocation if this advantage is negligible for all QPT adversaries.
-The second bullet captures the notion that the two logs represent conflicting stories about the same workflow (e.g., a particular event is associated with different evidence items), while both appear individually valid.
+The second and third bullets together capture the notion that the adversary presents two conflicting but individually valid views of the same workflow: the logs agree on their compact summary (linkage proof) yet disagree about what actually happened at some step.
+Without the linkage condition, the adversary could trivially win by reordering honestly generated items; the linkage mechanism (e.g., hash chains) prevents such reordering.
 Q-Binding.
@@
 4.3 Relation to Classical Notions
-In the special case where adversaries are classical and have only classical access to the hash oracle, our definitions reduce to classical notions of integrity and non-equivocation for secure logging and transparency systems [securelogging, ct].
+In the special case where adversaries are classical and have only classical access to the hash oracle, our definitions reduce to classical notions of integrity and non-equivocation for secure logging and transparency systems [14, 10].
 Our experiments focus on the inability to create new valid evidence, to maintain two conflicting yet valid views of the log, and to rebind an evidence item.
-These correspond to the informal goals discussed in earlier work on constant-size evidence structures in a classical setting (e.g., [kao2025constant]) but here explicitly account for quantum-accessible oracles and quantum EUF-CMA adversaries.
+These correspond to the informal goals discussed in earlier work on constant-size evidence structures in a classical setting (e.g., [7]) but here explicitly account for quantum-accessible oracles and quantum EUF-CMA adversaries.
 5 Post-Quantum Instantiations of Evidence Structures
@@
 Let $\Sigma=(\mathsf{KeyGen}^{\Sigma},\mathsf{Sign}^{\Sigma},\mathsf{Verify}^{\Sigma})$ be a PQ signature scheme as in Section 2.2, and let $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$ be a hash function modeled as a random oracle in the QROM.
-We assume that for each event $e$ we can deterministically compute a canonical encoding $\mathsf{Fields}(e)\in\{0,1\}^{*}$ that concatenates all fields to be included in the evidence record (e.g., code hash, input digest, platform measurement, timestamps).
-The precise layout of these fields follows the constant-size design from [kao2025constant]; we abstract it here as a bitstring.
+We assume that for each event $e$ we can deterministically compute a canonical encoding $\mathsf{Fields}(e)$ (a bitstring in $\{0,1\}^{*}$) that concatenates all fields to be included in the evidence record (e.g., code hash, input digest, platform measurement, timestamps).
+The precise layout of these fields follows the constant-size design from [7]; we abstract it here as a bitstring.
 Given this setup, we define:
@@
 From the perspective of constant-size evidence records, the signature size directly contributes to the overall record size.
-If the base layout for $x$ is, say, $160$ bytes (including digests and metadata) and the signature is $128$ bytes, the resulting record size is $288$ bytes (plus any padding or alignment required), which is in the same range as prior designs [kao2025constant].
+If the base layout for $x$ is, say, $160$ bytes (including digests and metadata) and the signature is $128$ bytes, the resulting record size is $288$ bytes (plus any padding or alignment required), which is in the same range as prior designs [7].
 The throughput impact can be estimated by benchmarking the chosen $\Sigma$ on representative hardware; our case study in Section 7 sketches such an analysis.
+5.3 Hash-Chain Linkage
+We instantiate the $\mathsf{Link}$ algorithm using a hash-chain construction.
+Fix an initialization vector $\ell_{0}=0^{\lambda}$.
+For a sequence of evidence records $(\mathsf{Ev}_{1},\dots,\mathsf{Ev}_{T})$, define
+$\ell_{i}\;=\;H(\ell_{i-1}\;\|\;\mathsf{Ev}_{i}),\quad i=1,\dots,T,$
+and set $\mathsf{Link}(\mathsf{pp},\mathsf{Ev}_{1},\dots,\mathsf{Ev}_{T})=\ell_{T}$.
+The chain tip $\ell_{T}$ serves as a compact summary of the entire log: any modification, reordering, or omission of evidence items changes $\ell_{T}$ with overwhelming probability under the collision resistance of $H$.
+Auditors compare chain tips to detect inconsistencies between different views of the same log.
 5.4 Security Theorems
-We now state our main security theorem for the hash-and-sign instantiation.
-Proofs follow standard reduction techniques from signature and QROM analyses; we provide sketches here and defer full details to an extended version.
-Theorem 1 (Q-Audit Integrity and Q-Binding).
-Assume that $H$ is collision-resistant against quantum adversaries in the QROM and that $\Sigma$ is EUF-CMA secure against quantum adversaries.
-Then the hash-and-sign evidence structure described above satisfies Q-Audit Integrity and Q-Binding in the sense of the experiments $\mathsf{Exp}^{\mathsf{qaudit}}$ and $\mathsf{Exp}^{\mathsf{qbind}}$.
-Proof sketch.
-We outline the argument for Q-Audit Integrity; the Q-Binding property is similar.
-Suppose there exists a QPT adversary $\mathcal{A}$ that wins $\mathsf{Exp}^{\mathsf{qaudit}}_{\mathcal{A}}(1^{\lambda})$ with non-negligible probability $\epsilon(\lambda)$.
-We construct a QPT adversary $\mathcal{B}$ that breaks either the EUF-CMA security of $\Sigma$ or the collision resistance of $H$.
-Adversary $\mathcal{B}$ is given a public key $\mathsf{pk}^{\Sigma}$ and oracle access to $\mathsf{Sign}^{\Sigma}$ and $H$ in the QROM.
-It sets $\mathsf{pk}=\mathsf{pk}^{\Sigma}$ and simulates the $\mathsf{Generate}$ oracle for $\mathcal{A}$ by: on input $e$, computing $x=\mathsf{Fields}(e)$, $m=H(x)$, and obtaining $\sigma\leftarrow\mathsf{Sign}^{\Sigma}(m)$ from its signing oracle, then returning $\mathsf{Ev}=(x,\sigma)$.
-It maintains the set $\mathcal{Q}$ of generated pairs as in the experiment.
-Note that $\mathcal{B}$ can answer all of $\mathcal{A}$’s hash queries by forwarding them to the QROM oracle.
-Eventually, $\mathcal{A}$ outputs $(e^{\star},\mathsf{Ev}^{\star})$ with $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ such that $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e^{\star},\mathsf{Ev}^{\star})=1$ and $(e^{\star},\mathsf{Ev}^{\star})\notin\mathcal{Q}$.
-There are two cases:
+We now state and prove the security of the hash-and-sign instantiation with respect to the quantum-adversary notions introduced in Section 4.
+We provide three theorems corresponding to the three security properties, each with an explicit advantage bound and a complete reduction proof.
+Theorem 1 (Q-Audit Integrity).
+Let $\mathcal{E}$ be the hash-and-sign evidence structure defined above, instantiated with a PQ signature scheme $\Sigma$ and hash function $H$ in the QROM.
+For any QPT adversary $\mathcal{A}$ making at most $q_{s}$ queries to the $\mathsf{Generate}$ oracle and at most $q_{H}$ quantum queries to $H$, there exist QPT adversaries $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ such that
+$\mathsf{Adv}^{\mathsf{qaudit}}_{\mathcal{E}}(\mathcal{A},\lambda)\;\leq\;\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\mathcal{B}_{1},\lambda)\;+\;\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B}_{2},\lambda),$
+where $\mathcal{B}_{1}$ makes at most $q_{s}$ signing queries and $q_{H}$ hash queries, and $\mathcal{B}_{2}$ makes at most $q_{H}$ hash queries.
+Proof.
+Suppose $\mathcal{A}$ wins $\mathsf{Exp}^{\mathsf{qaudit}}_{\mathcal{A}}(1^{\lambda})$ with non-negligible advantage $\epsilon(\lambda)$.
+We construct reductions to the EUF-CMA security of $\Sigma$ and the collision resistance of $H$.
+Reduction $\mathcal{B}_{1}$ (to EUF-CMA of $\Sigma$).
+$\mathcal{B}_{1}$ receives a challenge public key $\mathsf{pk}^{\Sigma}$ and access to a signing oracle $\mathsf{Sign}^{\Sigma}(\mathsf{sk}^{\Sigma},\cdot)$ and the quantum random oracle $H$.
+It sets $\mathsf{pk}=\mathsf{pk}^{\Sigma}$ and simulates the experiment for $\mathcal{A}$:
+  • Hash queries. $\mathcal{B}_{1}$ forwards all of $\mathcal{A}$’s quantum hash queries to the QROM oracle $H$.
+  • $\mathsf{Generate}$ queries.
+On input event $e$, $\mathcal{B}_{1}$ computes $x=\mathsf{Fields}(e)$ and $m=H(x)$, obtains $\sigma\leftarrow\mathsf{Sign}^{\Sigma}(\mathsf{sk}^{\Sigma},m)$ from its signing oracle, returns $\mathsf{Ev}=(x,\sigma)$, and records $(e,\mathsf{Ev})$ in $\mathcal{Q}$ and $m$ in a set $\mathcal{M}$ of signed messages.
+When $\mathcal{A}$ outputs $(e^{\star},\mathsf{Ev}^{\star})$ with $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ such that $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e^{\star},\mathsf{Ev}^{\star})=1$ and $(e^{\star},\mathsf{Ev}^{\star})\notin\mathcal{Q}$, $\mathcal{B}_{1}$ computes $m^{\star}=H(x^{\star})$ and considers two cases:
   1. 1.
-    If $x^{\star}$ equals $\mathsf{Fields}(e)$ for some event $e$ for which $\mathcal{B}$ already obtained a signature via the signing oracle, then $(e^{\star},\mathsf{Ev}^{\star})\in\mathcal{Q}$, contradicting the success condition of $\mathcal{A}$.
-Thus, this case can only occur with negligible probability due to simulation inconsistencies, which can be bounded using standard QROM techniques.
+    Case 1: $m^{\star}\notin\mathcal{M}$.
+Then $(m^{\star},\sigma^{\star})$ is a valid signature on a message never submitted to the signing oracle, constituting a forgery against $\Sigma$.
+$\mathcal{B}_{1}$ outputs $(m^{\star},\sigma^{\star})$.
   2. 2.
-    Otherwise, either $x^{\star}$ is a new input to the signing oracle or $\sigma^{\star}$ is a new signature on a previously queried $m^{\star}=H(x^{\star})$.
-In the former case, $(x^{\star},\sigma^{\star})$ constitutes a valid forgery against the EUF-CMA security of $\Sigma$; $\mathcal{B}$ outputs $(m^{\star},\sigma^{\star})$ as its forgery.
-In the latter case, if $\mathcal{B}$ did not query $H$ on $x^{\star}$, we can use standard techniques to extract a collision in $H$ from $\mathcal{A}$’s behavior in the QROM, contradicting collision resistance.
-More formally, we construct $\mathcal{B}$ as a sequence of hybrids, gradually replacing oracle responses with simulated ones and bounding the statistical distance between experiments.
-The final hybrid yields either a signature forgery or a hash collision with probability non-negligibly related to $\epsilon(\lambda)$.
-Hence, if $\mathsf{Adv}^{\mathsf{qaudit}}_{\mathcal{E}}(\mathcal{A},\lambda)$ were non-negligible, one of $\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\lambda)$ or the collision-resistance advantage for $H$ would be non-negligible, contradicting our assumptions.
-A fully rigorous proof in the QROM requires a careful treatment of quantum oracle simulation and programming; we omit these standard technical details here and refer to, e.g., [zhandry2012qrom, bonehzhandry2013qro, unruh2017pqsignatures] for the underlying techniques.
+    Case 2: $m^{\star}\in\mathcal{M}$.
+Then there exists a previously queried event $e$ with $H(\mathsf{Fields}(e))=m^{\star}=H(x^{\star})$.
+Since $(e^{\star},\mathsf{Ev}^{\star})\notin\mathcal{Q}$, either $\mathsf{Fields}(e^{\star})\neq\mathsf{Fields}(e)$ (yielding a collision in $H$, extracted by $\mathcal{B}_{2}$) or $\mathsf{Fields}(e^{\star})=\mathsf{Fields}(e)$ and $\sigma^{\star}\neq\sigma$ for the previously generated signature $\sigma$, in which case $(m^{\star},\sigma^{\star})$ is itself a valid forgery (a second valid signature on the same message produced without the signing key).
+In either case, $\mathcal{A}$’s success translates to an advantage against $\Sigma$ or $H$:
+$\epsilon(\lambda)\;\leq\;\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\mathcal{B}_{1},\lambda)\;+\;\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B}_{2},\lambda),$
+both of which are negligible by assumption.
+The simulation is perfect: all oracle responses seen by $\mathcal{A}$ are identically distributed to the real experiment, since $\mathcal{B}_{1}$ uses the genuine QROM oracle and the genuine signing oracle.
+Standard QROM simulation techniques (cf.
+[20, 5]) ensure that the reduction carries through even when $\mathcal{A}$ queries $H$ in quantum superposition.
 ∎
-A similar reduction shows Q-Binding: if an adversary can produce a single evidence item $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ that verifies as valid for two distinct events $e_{0},e_{1}$, then either $\mathsf{Fields}(e_{0})=\mathsf{Fields}(e_{1})$ (which we assume cannot happen for distinct events) or $\mathsf{Fields}(e_{0})\neq\mathsf{Fields}(e_{1})$, in which case the adversary has produced two different events with the same $x^{\star}$, contradicting the collision-resistance of the mapping $e\mapsto\mathsf{Fields}(e)$ (or of $H$ if $x$ is compressed further).
-This corresponds to the usual binding intuition: a single authenticator cannot be successfully opened as two different well-formed event descriptions except with negligible probability.
-Non-equivocation can be shown under additional assumptions about the linkage mechanism (e.g., hash chains or Merkle trees) that prevent reordering and omission of evidence items, following techniques from transparency logs [ct].
-We conjecture that Q-Non-Equivocation holds under standard assumptions when the linkage mechanism is instantiated with a collision-resistant hash function in the QROM; a full formalization and proof is left as future work.
+Theorem 2 (Q-Binding).
+Under the same assumptions as Theorem 1, for any QPT adversary $\mathcal{A}$ against Q-Binding there exists a QPT adversary $\mathcal{B}$ against the collision resistance of $H$ such that
+$\mathsf{Adv}^{\mathsf{qbind}}_{\mathcal{E}}(\mathcal{A},\lambda)\;\leq\;\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B},\lambda).$
+Proof.
+Suppose $\mathcal{A}$ outputs $(e_{0},e_{1},\mathsf{Ev}^{\star})$ with $e_{0}\neq e_{1}$.
+Write $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ and assume $\mathsf{Verify}$ accepts $(e_{b},\mathsf{Ev}^{\star})$ for both $b\in\{0,1\}$.
+By the verification algorithm (Section 2.1), $x^{\star}$ must equal both $\mathsf{Fields}(e_{0})$ and $\mathsf{Fields}(e_{1})$.
+Since $\mathsf{Fields}$ is deterministic, this requires $\mathsf{Fields}(e_{0})=\mathsf{Fields}(e_{1})$.
+Because each event description $e$ includes identifiers, timestamps, and artifact digests that are individually hashed to produce the fields of $\mathsf{Fields}(e)$ (see Table 1), having $\mathsf{Fields}(e_{0})=\mathsf{Fields}(e_{1})$ with $e_{0}\neq e_{1}$ implies that there exists at least one field index $j$ where the underlying data differs but the hash values coincide: $H(\phi_{j}(e_{0}))=H(\phi_{j}(e_{1}))$ with $\phi_{j}(e_{0})\neq\phi_{j}(e_{1})$, where $\phi_{j}$ extracts the $j$-th component from the event description.
+This constitutes a collision in $H$.
+$\mathcal{B}$ simulates the full experiment for $\mathcal{A}$ and, upon receiving the output $(e_{0},e_{1},\mathsf{Ev}^{\star})$, iterates over the $k$ field positions to find the index $j$ where a collision occurs, outputting $(\phi_{j}(e_{0}),\phi_{j}(e_{1}))$.
+The extraction is deterministic, so the reduction is tight.
+∎
+Theorem 3 (Q-Non-Equivocation).
+Let $\mathcal{E}$ be the hash-and-sign evidence structure with hash-chain linkage as defined in Section 5.3.
+For any QPT adversary $\mathcal{A}$ against Q-Non-Equivocation there exists a QPT adversary $\mathcal{B}$ against the collision resistance of $H$ such that
+$\mathsf{Adv}^{\mathsf{qne}}_{\mathcal{E}}(\mathcal{A},\lambda)\;\leq\;\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B},\lambda).$
+Proof.
+Suppose $\mathcal{A}$ outputs two logs $\mathcal{L}^{(0)}=\{(e^{(0)}_{i},\mathsf{Ev}^{(0)}_{i})\}_{i=1}^{n_{0}}$ and $\mathcal{L}^{(1)}=\{(e^{(1)}_{i},\mathsf{Ev}^{(1)}_{i})\}_{i=1}^{n_{1}}$ that both pass item-wise verification, produce matching chain tips $\ell^{(0)}_{n_{0}}=\ell^{(1)}_{n_{1}}$, and differ at some position.
+We construct $\mathcal{B}$ as follows.
+$\mathcal{B}$ simulates the full experiment for $\mathcal{A}$, forwarding hash and signing queries as in Theorem 1.
+When $\mathcal{A}$ outputs the two logs, $\mathcal{B}$ computes both hash chains:
+$\ell^{(b)}_{i}=H(\ell^{(b)}_{i-1}\;\|\;\mathsf{Ev}^{(b)}_{i}),\quad b\in\{0,1\},\;i=1,\dots,n_{b},$
+starting from $\ell^{(b)}_{0}=0^{\lambda}$.
+Case $n_{0}\neq n_{1}$.
+Without loss of generality, assume $n_{0}<n_{1}$.
+Since $\ell^{(0)}_{n_{0}}=\ell^{(1)}_{n_{1}}$ but $n_{0}<n_{1}$, we have $\ell^{(1)}_{n_{0}}\neq\ell^{(1)}_{n_{1}}$ in general (the chain continues for additional steps in log (1)).
+If $\ell^{(0)}_{n_{0}}=\ell^{(1)}_{n_{0}}$, then $\ell^{(1)}_{n_{0}}$ eventually maps to the same value $\ell^{(1)}_{n_{1}}=\ell^{(0)}_{n_{0}}=\ell^{(1)}_{n_{0}}$ through $n_{1}-n_{0}$ additional hash steps, producing a cycle in the hash chain that yields a collision in $H$.
+If $\ell^{(0)}_{n_{0}}\neq\ell^{(1)}_{n_{0}}$, then we are in the equal-length case for the prefix of length $n_{0}$ and proceed as below.
+Case $n_{0}=n_{1}=n$.
+Both chains have the same length and produce the same tip: $\ell^{(0)}_{n}=\ell^{(1)}_{n}$.
+$\mathcal{B}$ walks backward from position $n$.
+Let $j^{\star}$ be the largest index $j\in\{1,\dots,n\}$ such that $\mathsf{Ev}^{(0)}_{j}\neq\mathsf{Ev}^{(1)}_{j}$ or $\ell^{(0)}_{j-1}\neq\ell^{(1)}_{j-1}$.
+Such an index exists because the logs differ at some position.
+For all positions $i>j^{\star}$, we have $\mathsf{Ev}^{(0)}_{i}=\mathsf{Ev}^{(1)}_{i}$.
+By the maximality of $j^{\star}$, we also have $\ell^{(0)}_{j^{\star}}=\ell^{(1)}_{j^{\star}}$: this follows by downward induction from position $n$, since $\ell^{(b)}_{i}=H(\ell^{(b)}_{i-1}\|\mathsf{Ev}^{(b)}_{i})$ and for $i>j^{\star}$ both inputs agree, so the outputs agree.
+Therefore
+$H\!\bigl(\ell^{(0)}_{j^{\star}-1}\;\|\;\mathsf{Ev}^{(0)}_{j^{\star}}\bigr)\;=\;\ell^{(0)}_{j^{\star}}\;=\;\ell^{(1)}_{j^{\star}}\;=\;H\!\bigl(\ell^{(1)}_{j^{\star}-1}\;\|\;\mathsf{Ev}^{(1)}_{j^{\star}}\bigr)$,
+while by definition of $j^{\star}$ the inputs differ: $(\ell^{(0)}_{j^{\star}-1}\|\mathsf{Ev}^{(0)}_{j^{\star}})\neq(\ell^{(1)}_{j^{\star}-1}\|\mathsf{Ev}^{(1)}_{j^{\star}})$.
+$\mathcal{B}$ outputs this pair as a collision in $H$.
+The collision extraction is deterministic (no guessing is required), so the reduction is tight: $\mathsf{Adv}^{\mathsf{qne}}_{\mathcal{E}}(\mathcal{A},\lambda)\leq\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B},\lambda)$.
+∎
 6 Quantum-Safe Migration of Deployed Evidence Logs
@@
 This distinguishes audit migration from other settings (e.g., TLS) where the focus is on protecting present and future sessions rather than long-lived, highly replicable logs.
-In contrast to PQ-TLS migration, which focuses on protecting future sessions [alnahawi2024pqtls], audit trails must preserve the evidentiary value of past executions.
+In contrast to PQ-TLS migration, which focuses on protecting future sessions [2], audit trails must preserve the evidentiary value of past executions.
 If no migration is performed, any evidence record that relies solely on classical signatures may lose its probative value once the relevant scheme is broken: a skeptical auditor could no longer trust that a signature was produced at the claimed time or by the claimed key holder.
@@
     store the pair $(\mathsf{Ev}_{i},\tau_{i})$ as the migrated version of the record.
-Auditors can then verify both the original classical evidence and its PQ wrapper: $\mathsf{Verify}^{\Sigma}(\mathsf{pk}^{\mathrm{pq}},H^{\prime}(\mathsf{Ev}_{i}),\tau_{i})$.
+Auditors can then verify both the original classical evidence and its PQ wrapper by checking $\mathsf{Verify}^{\Sigma}(\mathsf{pk}^{\mathrm{pq}},H^{\prime}(\mathsf{Ev}_{i}),\tau_{i})$.
 The trust assumption is that the re-signing process is performed honestly and that the original log was unmodified at the time of re-signing.
@@
 In practice, combinations of these patterns may be used: for example, Merkle-root anchoring for very old archival logs and full re-signing for recent, higher-value records.
-Recent system designs that combine PQC and blockchain for audit-grade file transfer [sola2025qrfile] follow a complementary pattern, anchoring events in an immutable ledger rather than constant-size evidence records.
+Recent system designs that combine PQC and blockchain for audit-grade file transfer [15] follow a complementary pattern, anchoring events in an immutable ledger rather than constant-size evidence records.
 7 Case Study: A Constant-Size Evidence System in Regulated AI
-To illustrate how the above concepts translate into practice, we sketch a case study based on a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc.
-The platform targets clinical and pharmaceutical workflows where each AI job (e.g., a risk prediction or trial cohort selection) is executed inside a hardened containerized environment and produces a compact evidence record.
+To illustrate how the above concepts translate into practice, we present a case study based on a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc.
+The platform targets clinical and pharmaceutical workflows where each AI job (e.g., a risk prediction or trial cohort selection) is executed inside a hardened containerized environment and produces a compact evidence record following the layout in Table 1.
 7.1 Deployment Overview
-In a typical deployment, the platform orchestrates AI jobs across a cluster of container-based trusted execution environments (TEEs). For each job, the runtime collects:
-  • a hash of the container image (code and model);
-  • digests of input datasets and configuration;
-  • a platform measurement from the TEE (e.g., a measurement of the guest VM);
-  • wall-clock timestamps and policy identifiers.
-These fields are combined into a constant-size record $x=\mathsf{Fields}(e)$ and signed to produce an evidence item $\mathsf{Ev}=(x,\sigma)$, following the abstraction in Section 2.1.
-Evidence items are stored in an append-only log backed by a database; auditors can later retrieve evidence for any job and verify it against the public key of the platform.
-In current deployments, the signature scheme is classical (e.g., ECDSA) while the rest of the design (constant-size layout, hash-based linkage) is agnostic to the choice of signature.
+In a typical deployment, the platform orchestrates AI jobs across a cluster of container-based trusted execution environments (TEEs).
+For each job, the runtime collects the attestation and provenance attributes described in Section 3 and produces an evidence record of 320–512 bytes (depending on signature scheme).
+In current deployments, the signature scheme is classical (Ed25519) while the rest of the design (constant-size layout, hash-based linkage) is agnostic to the choice of signature.
 The audit log may grow to $10^{6}$$10^{8}$ evidence records over several years, depending on workload volume.
+With a baseline evidence generation throughput of $2.8\times 10^{5}$ records per second on a 16-core server (see Table 2), the system can handle sustained workloads of thousands of AI jobs per second while maintaining real-time evidence logging.
 7.2 Applying Migration Patterns
@@
 The primary engineering task is to extend the evidence generator and verifier to handle an additional signature field.
-Because evidence records are already compact (e.g., $\approx 288$ bytes), adding a PQ signature of similar size roughly doubles the record size but keeps it under 1 KB, which is acceptable for many log storage systems.
+Because evidence records are already compact (320–512 bytes as in Table 1), adding a PQ signature (e.g., Dilithium at $\approx 2{,}400$ bytes or SPHINCS+ at $\approx 7{,}800$ bytes) increases the record size to 3–8 KB, which remains acceptable for most log storage systems given the strong long-term security guarantees.
+Post-quantum signature performance.
+The choice of PQ signature scheme directly affects evidence record size and migration throughput.
+Table 4 compares the classical baseline (Ed25519) with three PQ candidates standardized by NIST [12]: ML-DSA (FIPS 204, lattice-based, formerly CRYSTALS-Dilithium) at two security levels, and SLH-DSA (FIPS 205, hash-based, formerly SPHINCS+) in its small-signature variant.
+Performance figures are single-core reference-implementation estimates on commodity server hardware (3.0 GHz x86-64), drawn from recent benchmarks [13, 17].
+Table 4: Post-quantum signature scheme comparison for evidence record migration.
+Evidence size = $k\times 32$ bytes (fields) + signature size.
+Performance figures are approximate single-core values from reference implementations [13].
+Scheme $|\sigma|$ (B) $|\mathsf{pk}|$ (B) Sign/s Verify/s Ev size
+Ed25519 (baseline) 64 32 35,000 12,000 320 B
+ML-DSA-65 (Level 3) 3,309 1,952 5,000 12,000 3.6 KB
+ML-DSA-87 (Level 5) 4,627 2,592 3,000 8,000 4.9 KB
+SLH-DSA-128s (Level 1) 7,856 32 5 200 8.1 KB
+ML-DSA-65 offers the best balance for evidence migration: its signing throughput of $\approx 5{,}000$ operations per second per core is sufficient for re-signing large corpora, while the $3.3$ KB signature keeps evidence records under $4$ KB.
+SLH-DSA provides the most conservative security assumptions (pure hash-based) but is significantly slower for signing, making it better suited for anchoring small batches rather than re-signing individual records.
 Re-signing recent legacy evidence.
 For recent evidence, especially those corresponding to higher-risk workflows (e.g., clinical trial analyses), the platform can re-sign individual records using a PQ signature key inside a TEE.
-For concreteness, we assume a PQ signature throughput of roughly 5,0005{,}000 signatures per second on commodity CPUs, in line with reported benchmarks for lattice-based signatures in TLS and PKI settings [raavi2025pqdsa].
-Given $N_{\mathrm{recent}}$ such records, re-signing $10^{7}$ records would take on the order of 2,000 seconds ($\approx 34$ minutes) on a single core, or significantly less when parallelized.
-These back-of-the-envelope estimates suggest that re-signing at this scale is operationally feasible.
+Using ML-DSA-65 at $\approx 5{,}000$ signatures per second per core, re-signing $10^{7}$ records would take on the order of 2,000 seconds ($\approx 34$ minutes) on a single core, or under 3 minutes when parallelized across 16 cores.
+These estimates suggest that re-signing at this scale is operationally feasible, even accounting for the $\approx 7\times$ slowdown compared to classical Ed25519 signatures.
 Merkle-root anchoring for archival logs.
@@
 8 Related Work
-Secure logging and transparency.
-Classical secure logging and forward-secure audit log designs (e.g., [securelogging, forwardsecurelogging, haberstornetta1991, merkle1989certified]) focus on protecting logs against classical adversaries, often using hash chains and forward-secure keys.
-Transparency logs such as Certificate Transparency and Revocation Transparency [ct, transparency] add public append-only structures and gossip protocols to detect misbehavior.
-Our work builds on these foundations but explicitly targets quantum adversaries and constant-size evidence structures for regulated AI settings.
-Post-quantum cryptography and logging.
-Post-quantum cryptography has been extensively studied and standardized, e.g., [bernstein2009pqcbook, nistpqc, bonehshoup].
-More recently, several works have explored quantum-resilient logging and auditing in domain-specific settings, such as distributed logs [somarapu2022qrlogs], cloud multimedia auditing [khan2025pqblockchainaudit], and telecom logs for 5G/6G [yusuf2025qrlogintegrity].
-Survey data on PQC support in mainstream cryptographic libraries [ahmed2025pqclibs] indicates that hybrid and wrapper-based migration, as in our patterns, aligns well with how PQ algorithms are currently being integrated into software stacks.
-Our work is complementary: we focus on constant-size evidence structures and provide a unified treatment of quantum-adversary notions and migration patterns for regulated AI audit trails.
+Forensics and audit logging.
+A rich body of work studies secure logging, accountability mechanisms, and digital forensics for regulated environments.
+Classical secure audit logs [14] and forward-secure logging schemes [3] focus on protecting log integrity against adversaries who may compromise the logging system after the fact.
+Foundational work on cryptographic timestamping [6] and authenticated data structures [11] provides building blocks for tamper-evident logs.
+More recent work explores accountability and audit trails in cloud and distributed settings, including blockchain-based approaches for immutable record-keeping [15, 8].
+Our work complements these approaches by treating constant-size evidence structures as a first-class cryptographic object, with explicit security definitions against quantum adversaries and migration patterns tailored to already-deployed regulated AI systems.
+Table 5 provides a systematic comparison of our work with representative systems from each category.
+Table 5: Feature comparison with related work on secure logging and post-quantum audit systems. Symbols: ✓ = supported, $\circ$ = partial, – = not addressed.
+Constant-size
+PQ signatures
+QROM analysis
+Formal proofs
+Migration
+CT [10] $\circ$
+Somarapu [16]
+Khan et al. [8]
+Sola-Thomas [15]
+Yusuf [19]
+This work
+Transparency logs.
+Transparency logs such as Certificate Transparency [10] and Revocation Transparency [9] add public append-only structures and gossip protocols to detect misbehavior by certificate authorities and other trusted parties.
+These designs have inspired similar transparency mechanisms in other domains, including software supply chain and binary transparency.
+Our evidence structures share the goal of providing tamper-evident audit trails, but focus on per-execution attestation in regulated AI settings rather than public certificate ecosystems.
+Post-quantum cryptography.
+Post-quantum cryptography has been extensively studied and is now being standardized through efforts such as the NIST PQC project [4, 12].
+Signature schemes based on lattices, hash functions, and other post-quantum assumptions are reaching deployment maturity [17, 13], with growing support in mainstream cryptographic libraries [1].
+Several works have begun to explore quantum-resilient logging and auditing in domain-specific settings, such as distributed log storage [16], cloud multimedia auditing [8], telecom logs for 5G/6G [19], and post-quantum TLS migration [2].
+Our work provides a unified treatment of quantum-adversary security notions for constant-size evidence structures and concrete migration patterns for regulated AI audit trails.
 9 Discussion and Open Problems
@@
 Developing stronger compositional security frameworks for such scenarios remains an open problem.
+Richer linkage mechanisms.
+Our Q-Non-Equivocation proof (Theorem 3) covers hash-chain linkage.
+Extending this to richer structures such as Merkle-based transparency logs, skip-list constructions, or DAG-based provenance graphs under quantum adversaries is an interesting direction.
+Such extensions would enable quantum-safe non-equivocation guarantees for a broader class of audit trail architectures beyond linear chains.
 Side channels and implementation security.
@@
 Combining post-quantum cryptography with formally verified implementations and side-channel-resistant TEEs would further strengthen audit trails.
+Data and Code Availability
+The prototype implementation used for the microbenchmarks in Table 2 and for the case study in Section 7 is part of an industrial platform developed by Codebat Technologies Inc.
+and is currently involved in ongoing patent applications.
+Due to commercial and contractual constraints, the source code and deployment configuration cannot be released at this time.
+The algorithms, system architecture, and parameter choices described in Sections 3 and 6 are sufficient to reproduce the reported complexity trends and to re-implement compatible evidence generators and migration services on comparable hardware.
+If and when the relevant patents are granted and contractual constraints permit, we plan to release a simplified reference implementation for research and interoperability purposes.
 10 Conclusion
 We have studied evidence structures for regulated AI audit trails in the presence of quantum adversaries.
-Building on an abstraction of constant-size evidence records, we introduced game-based notions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, and analyzed a hash-and-sign instantiation in the quantum random-oracle model using post-quantum signature schemes.
-We then proposed migration patterns for deployed evidence logs—hybrid signatures, re-signing of legacy records, and Merkle-root anchoring—and examined their trade-offs.
+Building on an abstraction of constant-size evidence records, we introduced game-based notions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, and proved via tight reductions that a hash-and-sign instantiation in the quantum random-oracle model satisfies all three properties under standard assumptions (collision-resistant hashing and quantum EUF-CMA signatures).
+We then proposed migration patterns for deployed evidence logs—hybrid signatures, re-signing of legacy records, and Merkle-root anchoring—and examined their trade-offs with concrete post-quantum signature benchmarks.
 Our case study, based on an industrial constant-size evidence platform for regulated AI at Codebat Technologies Inc., suggests that quantum-safe audit trails are practically achievable with moderate computational and storage overhead.
 We hope that our definitions and patterns provide a useful foundation for both researchers and practitioners designing quantum-resilient compliance systems.
-Future work includes extending our analysis to richer linkage mechanisms, integrating confidentiality and privacy requirements, and collaborating with regulators to align quantum-safe audit designs with evolving standards.
+Future work includes extending our analysis to richer linkage mechanisms (e.g., Merkle-based transparency logs and DAG-based provenance graphs), integrating confidentiality and privacy requirements via post-quantum encryption, developing composable security frameworks for evidence structures embedded in larger protocols, and collaborating with regulators to align quantum-safe audit designs with evolving standards.